40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 7 03:14:06 2013, Jenne, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
    Reply  Wed Aug 21 23:50:40 2013, Koji, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
       Reply  Fri Aug 23 01:52:34 2013, Jenne, Update, SUS, Violin filters moved to LSC, from SUS MICHloopNotches_23Aug2013.pdf
          Reply  Fri Aug 23 14:19:28 2013, rana, Update, SUS, Violin filters moved to LSC, from SUS 
Message ID: 8533     Entry time: Tue May 7 03:14:06 2013     Reply to this: 9048
Author: Jenne 
Type: Update 
Category: SUS 
Subject: PRM SUS_LSC violin (FM5) set to correct frequency 

While looking over Koji's shoulder earlier, I noticed the big peak in the PRM yaw spectrum (and I was starting to get annoyed by the hum....the fibox is so useful in motivating tasks that otherwise get looked over!) 

I used DTT's peak find feature (cursor tab, enable both cursors, select Peak X/Y as your 'statistic', set the 2 cursors to be on either side of the desired peak) to find the frequency of the PRM's violin mode.  It is 627.75 Hz. I adjusted FM5 of the C1:SUS-PRM_LSC filter bank (the "violin" filter) to be centered around this frequency, with the start and stop freqs +\- 4Hz.  I plotted the filter linearly in frequency to ensure that my target freq was not too close to either side of the bandstop.  After loading and engaging the new filter, the hum slowly started to go away. 

Note, for posterity:  The bandstop used to be centered around ~645 Hz or so.  I assume this is a copy-and-paste situation, where we hadn't gone through to check the exact frequency for each optic.

ELOG V3.1.3-